# ANID: How Far Are We? Evaluating the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v1

@@ 5. Human Assessment:
Human Validation
To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal. We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels. Participants rated images on a scale from 1 (poor) to 5 (excellent). Across all three dimensions, AI-generated images generally received lower scores, often between 2 and 3, compared to natural images.
Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts. The results revealed a noticeable gap between AIGC images and their natural counterparts, with the generated images often lacking the same level of detail and clarity.
Alignment: This aspect measured how well the generated images matched the given prompts, revealing a notable discrepancy. Generated images frequently deviated from the descriptions, indicating issues with model adherence to specified conditions.
Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality. While some generated images displayed impressive artistic qualities, the overall consistency and aesthetic appeal were generally considered inferior to that of natural images.
Overall, our human evaluation results are shown in Table 7 and Figure 5 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal. AIGC images consistently received lower scores, emphasizing the need for further advancements to narrow these gaps. These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
Moreover, the results from human validation indicate that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images. This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
More Evaluations
@@
To analyze the differences between AI-generated and natural images across various categories, we evaluated the performance of generated data from eight categories (’person’, ’animal’, ’indoor’, ’outdoor’, ’vehicle’, ’food’, ’sports’, ’accessory’) across three key aspects: (1) Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID); (2) Alignment, measured by CLIP Score; and (3) Aesthetic Quality, assessed by LAION-AES. Partial results are presented in Figures 6 and 7, while the full experimental findings are provided in Appendix Sec. E (Figures 13 and 14 ).
From Figure 6, we observe substantial variability in the differential rates for categories such as “food” and “sports”. Compared to “sports” images, the “food” demonstrate significantly higher discrepancies in image quality (CLIPIQA), revealing AI models struggle to generate natural food images. In Figure 7, the FID and SSIM metrics highlight distinct trends across categories. The “accessory” category exhibits the highest FID values, indicating poorer content distribution, while the “animal” category consistently achieves higher SSIM scores, reflecting better pixel-level similarity. These results suggest that certain categories, such as “accessory” and “indoor”, remain particularly challenging for AIGC models to reproduce accurately, especially in maintaining semantic consistency and achieving balanced content generation.
Overall, these findings demonstrate that AI-generated images still fall short of matching natural images in terms of naturalness and image quality. Additionally, the pronounced performance discrepancies across categories underscore an imbalance in the training data used for AIGC models. Addressing these discrepancies will require more balanced datasets and targeted improvements in model architecture and training strategies.
Figure 6: Differential rates of AI-generated images in the food and sports categories across naive image quality, alignment, and aesthetic metrics.
Figure 7: FID and SSIM metrics for AI-generated images across different source models and categories. FID evaluates distribution-level image quality, while SSIM measures pixel-level similarity to assess overall image quality.
Conclusion and Discussion
@@
In this paper, we develop a systematic and comprehensive framework to evaluate the discrepancies between AI-Natural images using the formulated DNAI dataset. Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability. While AIGC models have achieved impressive innovations, AIGIs still lag behind natural images in these critical areas. Besides, human evaluations also reveal lower scores for AIGC in terms of visual quality, semantic alignment, and aesthetics, emphasizing the gaps between AIGIs and natural images. Finally, the poor utility of AIGIs in downstream tasks indicates practical application challenges for AIGC.
@@
Appendix
Appendix Overview
@@
  • Section A: In-depth Description of the DNAI Dataset.
  • Section B: Detailed Overview of the Evaluated Generative Models.
@@
  • Section D: Extended Information on Human Perceptual Evaluations.
@@
A Details of the DNAI Dataset
Overview
@@ A Details of the DNAI Dataset
The DNAI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models. The natural images are sourced from the COCO validation set. AI-generated images were produced using eight different models. These images were generated under three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Representative samples from the DNAI dataset are shown in Figure 8. For a detailed breakdown of the DNAI dataset composition, please refer to Table 8.
Figure 8: Examples of AI-generated content (AIGC) created using five state-of-the-art generative models: SD_v1.4, SD_v1.5, Versatile Diffusion, SD_v2.1, and SD_XL. Each case is illustrated with three rows corresponding to Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I) generation modes. For the complete set of AIGC images, please refer to our DNAI dataset, which will be available publicly soon.
Types DFGAN GALIP SD_V1.4 SD_V1.5 Versartile SD_V2.1 SD_XL DALL·E 2
T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000
I2I - - 25,000 25,000 25,000 25,000 25,000 5,000
TI2I - - 25,000 25,000 25,000 25,000 25,000 -
Table 8: The details of the DNAI dataset.
Natural Images
@@
The DNAI dataset features a carefully curated selection of natural images sourced from the COCO validation set. This subset comprises 5,000 images, each paired with at least five unique captions, totaling 25,014 captions. These rich and varied textual descriptions enhance the dataset’s utility across a wide range of image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models. In this study, we use these images and their accompanying (five) captions to guide the generative models in creating new images.
Generated Images
@@
The DNAI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). These modes collectively contributed a significant volume of images to the dataset. The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, and DALL·E 2. Each model, except DALL·E 2, generated 25,000 images per guidance mode. DALL·E 2 produced 5,000 images per mode, resulting in a total of 440,000 generated images. It is important to note that DF-GAN and GALIP were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes. For detailed information, please refer to Table 8.
Image Resolution
The resolution of the generated images varies depending on the model used. Images generated by GALIP are 224x224 pixels, while those produced by DF-GAN are 256x256 pixels. The other models, including various versions of Stable Diffusion and DALL·E 2, generate images at a resolution of 512x512 pixels.
Dataset Composition
The DNAI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes. By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content. The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
B Evaluated Generative Models
@@
GALIP (Generative Adversarial Latent Image Processing): (Tao et al. 2023) is a generative model that produces high-quality images by leveraging adversarial techniques. It utilizes latent image processing to enhance both the quality and fidelity of generated content. GALIP’s architecture involves a generator and discriminator, where the generator creates images from latent vectors, and the discriminator evaluates their authenticity. Through iterative training, GALIP achieves highly detailed and realistic image synthesis.
@@
DFGAN (Deep Fusion Generative Adversarial Network): (Tao et al. 2022) focuses on generating images from text descriptions using deep fusion techniques. The model integrates textual information into the image generation process through multiple layers of fusion, ensuring alignment with input descriptions. DFGAN operates in two stages: the first generates a coarse image based on the text, and the second refines the image, enhancing detail and coherence.
@@
Stable Diffusion v1.4 (SD_V14): (Rombach et al. 2022) is part of the Stable Diffusion family, designed for image synthesis through iterative diffusion processes. Starting with pure noise, the model refines it into coherent images through a series of steps, guided by a learned model that predicts noise distribution. SD_V14 is recognized for producing high-resolution images with fine textures and complex structures.
@@
Stable Diffusion v1.5 (SD_V15): (Rombach et al. 2022) builds on SD_V14, introducing enhancements in architecture and training techniques to improve image quality, consistency, and the handling of complex scenes. While continuing to use the diffusion process, SD_V15 incorporates better optimization strategies and larger datasets, yielding superior results.
@@
Versatile Diffusion: (Xu et al. 2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution. Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
@@
Stable Diffusion v2.1 (SD_V21): (Rombach et al. 2022) is an advanced iteration of the Stable Diffusion series, offering significant improvements in image fidelity and generation speed. With optimizations in the diffusion process, better noise handling, and enhanced training algorithms, SD_V21 produces more realistic and high-quality images, making it highly effective for a range of creative and practical uses.
@@
Stable Diffusion XL (SD_XL): (Rombach et al. 2022) represents the most advanced model in the Stable Diffusion series, capable of generating extremely high-resolution images with intricate details. SD_XL uses an extended number of diffusion steps and a larger network architecture to manage increased complexity, making it ideal for applications requiring ultra-high resolution and precision, such as detailed artworks and large-scale prints.
@@
DALL·E 2: (Ramesh et al. 2022) is a generative model developed by OpenAI that excels in creating images from textual descriptions. As an advancement over the original DALL·E, it offers improved capabilities in generating high-quality, diverse, and coherent images. DALL·E 2 combines CLIP (Contrastive Language-Image Pre-training) with a transformer-based generative model to translate complex and abstract textual inputs into visual content, making it renowned for its creative and realistic outputs.
@@
Figure 9: We evaluate eight image-generating models from 13 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
C Evaluation Metrics
Overview
@@ C Evaluation Metrics
  • Pixel-level: Metrics such as PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity. These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
  • Frame-level: Metrics including PIQE, IL-NIQE, MUSIQ, DBCNN, CNNIQA, CLIPIQA, BRISQUE, TReS, HyperIQA, LIQE, and UNIQUE provide no-reference assessments of image quality. They evaluate attributes such as distortion levels, natural scene statistics, and learned features to predict the overall quality of the images. Additionally, naturalness metrics like NIQE and NRQM help gauge how closely AI-generated images resemble natural scenes based on statistical properties.
  • Visual Content Distribution: Metrics such as FID and Inception Score examine overall content distribution differences between AI-generated and natural images across the entire dataset. These metrics capture broad differences in content quality.
For alignment, we utilize the CLIP Score (Radford et al. 2021b) to measure the correspondence between text descriptions and the generated images. This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
In terms of aesthetic evaluation, we employ NIMA and LAION-AES metrics to assess the visual appeal and artistic quality of the images. These metrics predict human perception of image aesthetics based on deep learning models and various visual features.
@@
Naive Image Quality
@@
  • SSIM (Wang et al. 2004) (Structural Similarity Index) compares the structural information between a reference and a test image, considering luminance, contrast, and structure for a comprehensive measure of similarity.
@@
  • VIF (Sheikh and Bovik 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
  • VSI (Zhang, Shen, and Li 2014) (Visual Saliency-based Index) evaluates perceptual similarity by focusing on how salient (prominent) features in the images compare.
  • FSIM (Zhang et al. 2011) (Feature Similarity Index) measures similarity using phase congruency and gradient magnitude, emphasizing feature similarity, which is crucial for human perception.
  • LPIPS (Zhang et al. 2018) (Learned Perceptual Image Patch Similarity) uses deep network features to assess perceptual similarity, capturing human visual perception more effectively by comparing feature activations in a deep neural network.
  • DISTS (Ding et al. 2022) (Deep Image Structure and Texture Similarity) combines structural and textural similarity metrics using deep learning features to evaluate overall image similarity.
  • MAD (Larson and Chandler 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
@@
  • PIQE (N. et al. 2015) (Perception-based Image Quality Evaluator) provides a no-reference quality assessment by analyzing image blocks and estimating distortion levels.
  • IL-NIQE (Zhang, Zhang, and Bovik 2015) (Integrated Local Natural Image Quality Evaluator) evaluates image quality based on natural scene statistics and local features, offering a no-reference assessment.
  • MUSIQ (Ke et al. 2021) (Multi-Scale Image Quality) assesses image quality at multiple scales, considering both local and global features for a comprehensive no-reference evaluation.
  • DBCNN (Zhang et al. 2020) (Deep Bilinear Convolutional Neural Network) predicts image quality using deep learning, based on bilinear pooling of convolutional features.
  • LIQE (Ma et al. 2018) (Learning-based Image Quality Evaluator) combines traditional quality metrics with learned features for robust no-reference quality assessment.
  • CNNIQA (Kang et al. 2014) (Convolutional Neural Network Image Quality Assessment) leverages deep convolutional networks for no-reference image quality assessment using learned features.
  • CLIPIQA (Radford et al. 2021a) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
  • TReS (Radford et al. 2021a) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
  • HyperIQA (Su et al. 2020) (Hyper Network-based Image Quality Assessment) uses a hypernetwork to predict image quality, leveraging deep learning for no-reference quality assessment.
  • UNIQUE (Zhang et al. 2021) (Universal Quality Index with Deep Features) evaluates unique aspects of image quality using deep learning features, offering a no-reference assessment.
  • BRISQUE (Mittal, Moorthy, and Bovik 2012) (Blind/Referenceless Image Spatial Quality Evaluator) assesses image quality based on natural scene statistics, providing a no-reference measure.
  • NIQE (Mittal, Soundararajan, and Bovik 2013) (Natural Image Quality Evaluator) evaluates how natural an image appears by comparing its features with those derived from a dataset of natural images, without needing a reference image.
  • NRQM (Ma et al. 2018) (Naturalness Image Quality Metric) assesses how closely an image aligns with the statistical properties of natural scenes based on natural scene statistics and perceptual models.
@@
  • FID (Heusel et al. 2017) (Fréchet Inception Distance) measures the similarity between two sets of images by comparing the means and covariances of their feature representations, providing a comprehensive assessment of both quality and diversity.
  • Inception Score (Salimans et al. 2016) evaluates the quality of generated images by considering the confidence of the classifier and the diversity of the generated samples, with higher scores indicating better quality and diversity.
Alignment
To assess alignment, we employ the CLIP Score (Contrastive Language-Image Pre-training Score), which measures the correspondence between text descriptions and generated images. This metric evaluates how well the visual content aligns with the textual prompts by embedding both into a shared space and calculating their cosine similarity.
Aesthetic
  • NIMA (Esfandarani and Milanfar 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images. It generates a score that reflects human perception of visual appeal.
  • LAION-AES (Schuhmann et al. 2022) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.

### v2

@@ 4.6 Human Validation
4.7 Discrepancies Across Categories
@@
To analyze the differences between AI-generated and natural images across various categories, we evaluated the discrepancy of AIGI from eight categories (‘person’, ‘animal’, ‘indoor’, ‘outdoor’, ‘vehicle’, ‘food’, ‘sports’, ‘accessory’) across three key aspects: ➊ Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID, CLIP-FID, and Inception Score); ➋ Alignment, measured by CLIP Score; and ➌ Aesthetic Quality, assessed by LAION-AES. The results in Figure 8 reveal substantial variability across different categories, even for the AIGI from the same model (please refer to Appendix Sec. D.2, Figures 16 and 17 for integrated details). This could be attributed to the significant imbalance in the distribution of training data across these categories (Schuhmann et al., 2022b; Aghajanyan et al., 2023).
Figure 8: The difference rate on food and sport subcategories across quality, alignment, and aesthetic.
5 Conclusion and Discussion
@@
In this paper, we develop a systematic and comprehensive D-Judge to evaluate the discrepancies between AI-Natural images using the constructed D-ANI dataset. Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability, especially in the low-level and content distribution-level image quality and fine-grained downstream tasks, like object recognition and object VAQ. While AIGC models have achieved impressive innovations, AIGIs still lag behind natural images in these critical areas. Besides, human evaluations also reveal lower scores for AIGC in terms of visual quality, semantic alignment, and aesthetics, emphasizing the gaps between AIGIs and natural images. Finally, the poor utility of AIGIs in downstream tasks indicates practical application challenges for AIGC.
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix Overview
@@
  • Section B: In-depth Description of the D-ANI Dataset.
@@
  • Section D: Additional Evaluation Results.
@@
Appendix B Details of the D-ANI Dataset
B.1 Overview
@@
The D-ANI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models. The natural images are sourced from the COCO validation set. AI-generated images were produced using eight different models. These images were generated under three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Representative samples from the D-ANI dataset are shown in Figures 9,  10 and  11. For a detailed breakdown of the D-ANI dataset composition, please refer to Table 6.
Figure 9: Examples of AI-generated content (AIGC) created using Text-to-Image (T2I) guidance.
Figure 10: Examples of AI-generated content (AIGC) created using Image-to-Image (I2I) guidance.
Figure 11: Examples of AI-generated content (AIGC) created using Text-and-Image-to-Image (TI2I) guidance.
B.2 Natural Images
@@
The D-ANI dataset features a carefully curated selection of natural images sourced from the COCO validation set. This subset comprises 5,000 images, each paired with at least five unique captions, totaling 25,014 captions. These rich and varied textual descriptions enhance the dataset’s utility across a wide range of image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models. In this study, we use these images and their accompanying (five) captions to guide the generative models in creating new images.
B.3 Generated Images
@@
The D-ANI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). These modes collectively contributed a significant volume of images to the dataset. The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, DALL·E 2 and DALL·E 3. Each model, except DALL·E 2 and DALL·E 3, generated 25,000 images per guidance mode. DALL·E 2 and DALL·E 3 produced 5,000 images for per mode, resulting in a total of 440,000 generated images. It is important to note that DF-GAN, GALIP and DALL·E 3 were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes. For detailed information, please refer to Table 6.
The involved generative models are as follows:
@@
  • GALIP (Generative Adversarial Latent Image Processing): (Tao et al., 2023) is a generative model that produces high-quality images by leveraging adversarial techniques. It utilizes latent image processing to enhance both the quality and fidelity of generated content. GALIP’s architecture involves a generator and discriminator, where the generator creates images from latent vectors, and the discriminator evaluates their authenticity. Through iterative training, GALIP achieves highly detailed and realistic image synthesis.
@@
  • DFGAN (Deep Fusion Generative Adversarial Network): (Tao et al., 2022) focuses on generating images from text descriptions using deep fusion techniques. The model integrates textual information into the image generation process through multiple layers of fusion, ensuring alignment with input descriptions. DFGAN operates in two stages: the first generates a coarse image based on the text, and the second refines the image, enhancing detail and coherence.
@@
  • Stable Diffusion v1.4 (SD14): (Rombach et al., 2022) is part of the Stable Diffusion family, designed for image synthesis through iterative diffusion processes. Starting with pure noise, the model refines it into coherent images through a series of steps, guided by a learned model that predicts noise distribution. SD14 is recognized for producing high-resolution images with fine textures and complex structures.
@@
  • Stable Diffusion v1.5 (SD15): (Rombach et al., 2022) builds on SD14, introducing enhancements in architecture and training techniques to improve image quality, consistency, and the handling of complex scenes. While continuing to use the diffusion process, SD15 incorporates better optimization strategies and larger datasets, yielding superior results.
@@
  • Versatile Diffusion: (Xu et al., 2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution. Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
@@
  • Stable Diffusion v2.1 (SD21): (Rombach et al., 2022) is an advanced iteration of the Stable Diffusion series, offering significant improvements in image fidelity and generation speed. With optimizations in the diffusion process, better noise handling, and enhanced training algorithms, SD21 produces more realistic and high-quality images, making it highly effective for a range of creative and practical uses.
@@
  • Stable Diffusion XL (SDxl): (Rombach et al., 2022) represents the most advanced model in the Stable Diffusion series, capable of generating extremely high-resolution images with intricate details. SDxl uses an extended number of diffusion steps and a larger network architecture to manage increased complexity, making it ideal for applications requiring ultra-high resolution and precision, such as detailed artworks and large-scale prints.
@@
  • DALL·E 2 (D·E 2): (Ramesh et al., 2022) is a generative model by OpenAI designed to create images from textual descriptions. It produces high-quality, diverse, and coherent outputs by combining CLIP (Contrastive Language-Image Pre-training) with a transformer-based generative architecture. DALL·E 2 excels at translating complex and abstract textual prompts into visually captivating content, establishing itself as a benchmark model in text-to-image generation for its creativity and realism.
  • DALL·E 3 (D·E 3): (OpenAI, 2024) is the latest evolution in OpenAI’s text-to-image generative models, offering significant advancements over DALL·E 2. It excels in interpreting nuanced and detailed prompts, producing outputs that are more coherent, visually appealing, and aligned with the given descriptions. DALL·E 3 integrates seamlessly with GPT-4, leveraging its advanced language understanding to ensure a deeper alignment between textual input and generated images. This integration enhances both the quality and controllability of image generation, making DALL·E 3 a powerful tool for a wide range of creative and professional applications.
@@
B.4 Image Resolution
The resolution of the generated images varies across models. GALIP produces images at 224x224 pixels, while DF-GAN generates slightly higher-resolution images at 256x256 pixels. Models such as various versions of Stable Diffusion and DALL·E 2 generate images at a resolution of 512x512 pixels. Notably, DALL·E 3 delivers the highest resolution among these models, producing images at 1024x1024 pixels.
B.5 Composition
The D-ANI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes. By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content. The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
Table 6: The details of the D-ANI dataset.
Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000 5,000
I2I - - 25,000 25,000 25,000 25,000 25,000 5,000 -
TI2I - - 25,000 25,000 25,000 25,000 25,000 - -
Appendix C Evaluation Metrics
C.1 Overview
@@
  • Pixel-level: Metrics such as SSIM, PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity. These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
  • Frame-level: Metrics including PIQE, IL-NIQE, MUSIQ, DBCNN, CNNIQA, CLIPIQA, BRISQUE, TReS, HyperIQA, LIQE, UNIQUE and QAlign provide no-reference assessments of image quality. They evaluate attributes such as distortion levels, natural scene statistics, and learned features to predict the overall quality of the images. Additionally, naturalness metrics like NIQE and NRQM help gauge how closely AI-generated images resemble natural scenes based on statistical properties.
  • Visual Content Distribution: Metrics such as FID, CLIP-FID and Inception Score examine overall content distribution differences between AI-generated and natural images across the entire dataset. These metrics capture broad differences in content quality.
For alignment, we utilize the CLIP Score (Radford et al., 2021a) to measure the correspondence between text descriptions and the generated images. This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
In terms of aesthetic evaluation, we employ NIMA, LAION-AES and QAlign metrics to assess the visual appeal and artistic quality of the images. These metrics predict human perception of image aesthetics based on deep learning models and various visual features.
@@
C.2 Naive Image Quality
@@
  • SSIM (Wang et al., 2004) (Structural Similarity Index) compares the structural information between a reference and a test image, considering luminance, contrast, and structure for a comprehensive measure of similarity.
@@
  • VIF (Sheikh & Bovik, 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
  • VSI (Zhang et al., 2014) (Visual Saliency-based Index) evaluates perceptual similarity by focusing on how salient (prominent) features in the images compare.
  • FSIM (Zhang et al., 2011) (Feature Similarity Index) measures similarity using phase congruency and gradient magnitude, emphasizing feature similarity, which is crucial for human perception.
  • LPIPS (Zhang et al., 2018) (Learned Perceptual Image Patch Similarity) uses deep network features to assess perceptual similarity, capturing human visual perception more effectively by comparing feature activations in a deep neural network.
  • DISTS (Ding et al., 2022) (Deep Image Structure and Texture Similarity) combines structural and textural similarity metrics using deep learning features to evaluate overall image similarity.
  • MAD (Larson & Chandler, 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
@@
  • PIQE (N. et al., 2015) (Perception-based Image Quality Evaluator) provides a no-reference quality assessment by analyzing image blocks and estimating distortion levels.
  • IL-NIQE (Zhang et al., 2015) (Integrated Local Natural Image Quality Evaluator) evaluates image quality based on natural scene statistics and local features, offering a no-reference assessment.
  • MUSIQ (Ke et al., 2021) (Multi-Scale Image Quality) assesses image quality at multiple scales, considering both local and global features for a comprehensive no-reference evaluation.
  • DBCNN (Zhang et al., 2020) (Deep Bilinear Convolutional Neural Network) predicts image quality using deep learning, based on bilinear pooling of convolutional features.
  • LIQE (Ma et al., 2018) (Learning-based Image Quality Evaluator) combines traditional quality metrics with learned features for robust no-reference quality assessment.
  • CNNIQA (Kang et al., 2014) (Convolutional Neural Network Image Quality Assessment) leverages deep convolutional networks for no-reference image quality assessment using learned features.
  • CLIPIQA (Radford et al., 2021b) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
  • TReS (Radford et al., 2021b) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
  • HyperIQA (Su et al., 2020) (Hyper Network-based Image Quality Assessment) uses a hypernetwork to predict image quality, leveraging deep learning for no-reference quality assessment.
  • UNIQUE (Zhang et al., 2021) (Universal Quality Index with Deep Features) evaluates unique aspects of image quality using deep learning features, offering a no-reference assessment.
  • BRISQUE (Mittal et al., 2012) (Blind/Referenceless Image Spatial Quality Evaluator) assesses image quality based on natural scene statistics, providing a no-reference measure.
  • NIQE (Mittal et al., 2013) (Natural Image Quality Evaluator) evaluates how natural an image appears by comparing its features with those derived from a dataset of natural images, without needing a reference image.
  • NRQM (Ma et al., 2018) (Naturalness Image Quality Metric) assesses how closely an image aligns with the statistical properties of natural scenes based on natural scene statistics and perceptual models.
  • QAlign (Wu et al., 2024b) evaluates image quality by assessing its alignment with semantic and perceptual features indicative of high visual fidelity.
@@
  • FID (Heusel et al., 2017) (Fréchet Inception Distance) measures the similarity between two sets of images by comparing the means and covariances of their feature representations, providing a comprehensive assessment of both quality and diversity.
  • CLIP-FID (Kynkäänniemi et al., 2023) (CLIP Fréchet Inception Distance) is an extension of FID that measures the similarity between two sets of images by utilizing CLIP’s multi-modal feature space. Instead of relying solely on Inception-based features, CLIP-FID leverages the semantic-rich embeddings from CLIP, which are aligned across image and text modalities. This provides a more nuanced assessment of image quality and semantic alignment, particularly for generative models designed to capture complex textual descriptions.
  • Inception Score (Salimans et al., 2016) evaluates the quality of generated images by considering the confidence of the classifier and the diversity of the generated samples, with higher scores indicating better quality and diversity.
C.3 Alignment
To assess alignment, we employ the CLIP Score (Contrastive Language-Image Pre-training Score, CS), which measures the correspondence between text descriptions and generated images. This metric evaluates how well the visual content aligns with the textual prompts by embedding both into a shared space and calculating their cosine similarity.
C.4 Aesthetic
  • NIMA (Esfandarani & Milanfar, 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images. It generates a score that reflects human perception of visual appeal.
  • LAION-AES (Schuhmann et al., 2022a) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
  • QAlign (Wu et al., 2024b) assesses image aesthetics by analyzing its adherence to human visual preferences and artistic principles, leveraging semantic coherence and appeal.